t-7 divided by 4 then plus 5 =16

Answer by checkley79(3341)

(t-7)/4+5=16
(t-7)/4=16-5
(t-7)/4=11
t-7=11*4
t-7=44
t=44+7
t=51 ans.
proof:
(51-7)/4+5=16
44/4=16-5
11=11